# Generative Inverse Design of Metamaterials with Functional Responses by Interpretable Learning

### v3

@@ 3.3 Synthetic Design Problems for Rapid Validation and Visualization
Figure 6: Visualization of estimated likelihood and validation metrics for synthetic problems. (A) Likelihood function values for randomly created design targets. Orange lines show boundaries of actual feasible regions associated with the targets $\mathcal{T}=\{I(a,b;s)=1|\forall s\in\Omega\}$. (B) Likelihood function values estimated by random forests with varying numbers of decision trees. The design target is set as $\mathcal{T}=\{I(a,b;s)=1|\forall s\in[0.45,0.48]\}$ for the SqExp problem and $\mathcal{T}=\{I(a,b;s)=1|\forall s\in[0.63,0.68]\cup[0.69,0.71]\}$ for the SupSin problem. (C) Validation metrics for inverse design generation.
For each synthetic example, we split the data into 80 training data and 20 test data, and trained a random forest, with the same hyperparameter settings as the other problems, to predict the binary indicators $I(a,b;s)$. The F1 scores are 0.85 and 0.86 for the SqExp and the SupSin problems, respectively. The resulting confusion matrices are shown in SI Appendix, Tables S3-S4. We evaluated the inverse design performance with the trained models.
@@ 4 Summary and Outlook
RIGID offers two levels of interpretability. Firstly, RIGID is more transparent than deep learning models. The transparency of the random forest as a forward model helps us understand how the predictions are made. The inverse design process based on likelihood estimation and sampling is also transparent. The transparency of both forward and inverse processes offers insights into the decision-making process and diagnostic capabilities. Secondly, owing to the interpretability of the forward prediction model, we can derive the likelihood, which the inverse design relies on, by probing into each decision path of trees in the random forest. The likelihood is an interpretable metric to help us understand the probability of a generated design satisfying the given target and thus aid the decision-making process when selecting final solutions. The results (Figures 3D, 3E, 4D, 4E, and 6C) have demonstrated a high correlation between the likelihood and actual probability of target satisfaction. The likelihood also serves as an indicator of the need for adjusting the design space. When the likelihood is low across the entire design space, the model believes the target is almost impossible to achieve, which may indicate a need for expanding the design space (i.e., extending design space bounds or adding new design freedom) to allow wider coverage of the response space and easier discovery of satisfying solutions. Please see more discussion in SI Appendix, Sec. D. Another benefit of having the explicit likelihood as an interpretable metric is that we can adjust the likelihood threshold to trade-off exploration and exploitation — higher (lower) threshold means more focus on exploitation (exploration). Both levels of interpretability were absent from most prior works of iteration-free, single-shot inverse design methods for metamaterials.
This work focuses on the scenario of small data, low-dimensional problems. As the problem dimension increases, the requirement for dataset size will quickly increase due to the curse of dimensionality. While there is a large collection of deep learning methods to address high-dimensional and large dataset problems, the setting of small data and low dimensionality seems to be under-studied by prior works, despite this setting’s practicality under constrained budgets. RIGID can fill this gap.

### v4

@@ 3.3 Synthetic Design Problems for Rapid Validation and Visualization
Figure 7: Visualization of estimated likelihood and validation metrics for synthetic problems. (A) Likelihood function values for randomly created design targets. Orange lines show boundaries of actual feasible regions associated with the targets $\mathcal{T}=\{I(a,b;s)=1|\forall s\in\Omega\}$. Points show satisfactory RIGID designs and GA designs. (B) Likelihood function values estimated by random forests with varying numbers of decision trees. The design target is set as $\mathcal{T}=\{I(a,b;s)=1|\forall s\in[0.45,0.48]\}$ for the SqExp problem and $\mathcal{T}=\{I(a,b;s)=1|\forall s\in[0.63,0.68]\cup[0.69,0.71]\}$ for the SupSin problem. (C) Validation metrics for inverse design generation.
For each synthetic example, we split the data into 80 training data and 20 test data. We trained a random forest with the same hyperparameter settings as the other problems, to predict the binary indicators $I(a,b;s)$. The F1 scores are 0.85 and 0.86 for the SqExp and the SupSin problems, respectively. The resulting confusion matrices are shown in SI Appendix, Tables S3-S4. We evaluated the inverse design performance with the trained models.
@@
We also visualize the distribution of satisfactory designs generated by both RIGID and GA in Fig. 7A. Again, we observe a much broader coverage of the feasible design space by RIGID designs, while GA designs are merely minor perturbations of one or two feasible solutions.
@@
4 Discussion
In this section, we discuss the benefits and necessities of RIGID in terms of its interpretability, data demand, the ability to deal with discrete design variables, and the diversity of generated designs.
4.1 Interpretability
RIGID offers two levels of interpretability. Firstly, RIGID is more transparent than deep learning models. The transparency of the random forest as a forward model helps us understand how the predictions are made. The inverse design process based on likelihood estimation and sampling is also transparent. The transparency of both forward and inverse processes offers insights into the decision-making process and diagnostic capabilities. Secondly, owing to the interpretability of the forward prediction model, we can derive the likelihood, which the inverse design relies on, by probing into each decision path of trees in the random forest. The likelihood is an interpretable metric to help us understand the probability of a generated design satisfying the given target and thus aid the decision-making process when selecting final solutions. The results (Figures 3D, 3E, 5D, 5E, and 7C) have demonstrated a high correlation between the likelihood and actual probability of target satisfaction. The likelihood also serves as an indicator of the need for adjusting the design space. When the likelihood is low across the entire design space, the model believes the target is almost impossible to achieve, which may indicate a need for expanding the design space (i.e., extending design space bounds or adding new design freedom) to allow wider coverage of the response space and easier discovery of satisfying solutions. Please see more discussion in SI Appendix, Sec. S4. Another benefit of having the explicit likelihood as an interpretable metric is that we can adjust the likelihood threshold to trade-off exploration and exploitation — higher (lower) threshold means more focus on exploitation (exploration). Both levels of interpretability were absent from most prior works of iteration-free, single-shot inverse design methods for metamaterials.
4.2 Data Demand
This work focuses on the scenario of small data, low-dimensional problems. As the problem dimension increases, the requirement for dataset size will quickly increase due to the curse of dimensionality. While there is a large collection of deep learning methods to address high-dimensional and large dataset problems, the setting of small data and low dimensionality seems to be under-studied by prior works, despite this setting’s practicality under constrained budgets. RIGID fills this gap by offering a much more data-efficient and interpretable alternative to deep-learning-based single-shot inverse design methods.
4.3 Diversity versus Feasibility of Generated Designs
For all four problems demonstrated in this work, when the sampling threshold is set to 0.8, RIGID achieves a satisfaction rate of 100%. Although GA can also lead to a high target satisfaction rate for relatively simple inverse design problems (e.g., the optical metasurface design problem in Sec. 3.2), the generated solutions lack diversity. In contrast, RIGID provides a unique way to control the trade-off between the diversity and the feasibility of generated designs by tuning the sampling threshold. It is important to note that, without considering diversity, a high target satisfaction rate becomes trivial, as we can simply apply small perturbations to one feasible solution to generate an infinite number of similar solutions that have similar responses and hence are also likely to meet the target. In many inverse design problems where solutions are non-unique, finding one solution can be simple, but the real challenge lies in discovering a broad range of feasible solutions—a problem where RIGID will outperform optimization-based methods as demonstrated by our results.
4.4 Discrete Design Variables